# CoRE: LLM as Interpreter for Natural Language Programming, Pseudo-Code Programming, and Flow Programming of AI Agents

### v1

@@ Abstract
Since their inception, programming languages have trended towards greater readability and lower barriers for programmers. Following this trend, natural language can be a promising type of programming language that provides great flexibility and usability and helps towards the democracy of programming. However, the inherent vagueness, ambiguity, and verbosity of natural language pose significant challenges in developing an interpreter that can accurately understand the programming logic and execute instructions written in natural language. Fortunately, recent advancements in Large Language Models (LLMs) have demonstrated remarkable proficiency in interpreting complex natural language. Inspired by this, we develop a novel system for Code Representation and Execution (CoRE), which employs LLM as interpreter to interpret and execute natural language instructions. The proposed system unifies natural language programming, pseudo-code programming, and flow programming under the same representation for constructing language agents, while LLM serves as the interpreter to interpret and execute the agent programs. In this paper, we begin with defining the programming syntax that structures natural language instructions logically. During the execution, we incorporate external memory to minimize redundancy. Furthermore, we equip the designed interpreter with the capability to invoke external tools, compensating for the limitations of LLM in specialized domains or when accessing real-time information. This work is open-source at https://github.com/agiresearch/CoRE.
@@ 2.2 Large Language Models and AI Agents for Problem Solving
Large Language Models (LLMs) have emerged as powerful tools for problem solving, encompassing tasks in reasoning, planning, and code generation. LLM reasoning typically involves decomposing a complex task into a sequence of steps, also known as a reasoning chain [48]. Prominent approaches in LLM reasoning include Chain-of-Thought (CoT) and its derivatives [48, 25]. To further improve the reasoning ability of LLM, several work has been proposed. The Self-consistency method [46] samples multiple reasoning paths and selects the most consistent outcome by voting. Additionally, classical data structures like trees and graphs are utilized to enhance reasoning efficiency and accuracy in fewer steps [51, 3]. Apart from reasoning, planning is also an important task that can be used to solve problems. LLM Planning involves generating a series of actions to achieve the predefined goals [16]. Recent advancements include direct prompting of LLMs for planning tasks, showing promising results [20, 44, 11]. Finite state machines have been integrated into LLM to enhance the planning ability [28, 50]. ReAct [52] proposes to leverage external tools like search engine to enhance the LLM planning. Besides, considering the powerful ability of LLM in programming, recent work propose to generate programming code to solve problems [31, 22, 30, 6, 39, 35]. Furthermore, the “self-reflection” mechanism [32, 38, 43] enables LLMs to critique their own outputs, significantly enhancing performance in tasks such as reasoning [3] and code generation [7]. In contrast to existing methods that directly use LLMs for generating solutions, the proposed CoRE system utilizes LLMs as interpreters, executing solutions designed by humans to address complex questions. This approach leverages human creativity in solution design, coupled with LLM’s ability, to enhance problem-solving capabilities in natural language programming contexts.
@@ 3.1 CoRE Language Syntax
Programming requires logical thinking and the ability to break down complex problems into smaller, more manageable instructions. To organize natural language instructions in logical thinking, we define the basic structural representation for each step, which consists of four components. An example can be found in Figure 1.
  • Step Name: Each step in the program is uniquely identified by a step name. This identifier serves a similar functionality to function identifiers in traditional programming languages, facilitating navigation and reference within the program structure. It ensures that each operation within the program can be distinctly addressed and accessed.
  • Step Type: The step type categorizes the nature of the operation being performed in each step, analogous to control structures in conventional programming. We define three primary types of steps:
@@
    • Decision: Corresponding to conditional statements (e.g., ’if-else’), this step involves branching the program flow based on evaluated conditions, leading to multiple potential paths.
    • Terminal: Similar to the ’end’ or ’return’ statement, this step marks the conclusion of the program, indicating that no further steps are to be executed.
  • Step Instruction: The step instruction explicates the task to be conducted at a particular step. This component is integral as it provides the semantic content necessary for execution, paralleling the statement block in traditional programming languages. By demonstrating operations in natural language, NLPg lowers the barrier to programming, making it more readable for non-expert programmers.
  • Step Connection: Step connections define the progression from one step to another, establishing the flow the program execution. In process steps, a single subsequent step is specified. In decision steps, multiple pathways are delineated based on conditions. Terminal steps, by definition, do not lead to any future steps, indicating the end of program execution.
For each step in the program, the above four components are separated by ‘:::’ (as illustrated in the CoRE language in Figure 1) since it is unusual in natural language instructions. It is very flexible to use other special signals to separate different components.
In programming languages, there are three basic control constructs in programming [9, 40]: sequence, selection, and iteration. These three basic constructs can be easily designed within our CoRE language.
  • Sequence: Sequence in programming is the execution of statements in a linear order, with each statement leading to the next. In the NLPg framework, this construct is designed by setting the “Step Connection” to point to the subsequent step. Each step operates under the Process type until the sequence concludes.
@@
  • Iteration: Iteration involves repeating a set of operations until a certain condition is met, akin to loops in conventional programming. In our CoRE language, we utilize a step with the Decision type to assess whether the loop condition has been fulfilled. At the end of one loop cycle, the “Step Connection” is configured to point back to the previous Decision step, thereby enabling the continuation of the loop.
@@ 3.2 LLM as Interpreter
In this section, we will discuss how the CoRE system utilizes a Large Language Model (LLM) as an interpreter to execute programs written in the CoRE language. We will demonstrate the execution of a single step within the CoRE system, which is illustrated in Figure 3. More specifically, the system executes a single step in four procedures. First of all, the interpreter determines the useful information to execute the current step. Then the interpreter will integrate all relevant information to construct the prompt. Based on the generated prompt, the interpreter will generate response and may utilize tools to execute the current step. Finally, after executing the current step, the interpreter will determine the next step based on step type and execution results. We will introduce four parts in details.
@@ 3.2.1 Observation Retrieval from Memory
This initial procedure is critical because it sets the stage for the entire execution process of the current step. The system’s memory serves as a repository of all prior observations related to the program, where the observation represents the results of tool execution, such as search results. During this phase, the interpreter scans the memory to identify records that are relevant to the current instruction. This selective retrieval ensures that the interpreter’s decisions are informed by accurate and contextually relevant data, which is crucial for the correctness and efficiency of the program. We present an example in Figure 4.
@@ 3.2.2 Input Prompt Construction
Constructing the prompt is essentially about synthesizing information into a comprehensive and coherent query that the LLM can understand and respond to effectively. This involves combining multiple information into a single, structured prompt that guides the LLM towards generating the most appropriate and contextually relevant response. In our CoRE system, the interpreter constructs a detailed prompt with four elements:
@@ 3.2.3 Output Analysis
While the LLM can generate direct responses, complex tasks might require capabilities beyond its immediate scope. Incorporating the use of specialized tools when necessary extends the LLM’s capabilities, allowing the system to handle a broader range of tasks effectively. In our CoRE system, the interpreter will make a decision about whether employ specialized tools based on the LLM’s initial response and the demands of the task at hand, which ensures that the system remains highly functional and versatile, actively solving problems rather than merely processing language for the current step. Specifically, if tool usage is warranted, the system will select the suitable tool, configure it with the necessary parameters, execute it, and integrate the output into the ongoing process. We demonstrate the execution process with an example shown in Figure 5.
3.2.4 Logic Representation
@@
Determining the appropriate next step in the program is critical, especially in multi-path scenarios where different outcomes can lead to different subsequent actions. In the CoRE language, the Decision steps indicate multiple path with corresponding conditions. The system intelligently interprets the LLM’s responses — even those that do not exactly match predefined conditions—using semantic analysis. This adaptive approach allows the system to navigate through decision points effectively, ensuring logical progression toward the program’s goals. An example has been shown in Figure 6.
@@ 4.1 Backbone Large Language Model (LLM)
We conduct the experiments of our proposed system CoRE on two representative LLMs as interpreters:
  • GPT-4 [36] (Closed-source) is a generative pre-trained transformer of OpenAI. In this work, we use the “GPT-4-1106-preview” version.
@@ 4.2 Planning Schema of LLMs
  • Chain-of-Thought (CoT) [48] induces the LLM to generate a coherent language sequence that serves as a meaningful intermediate step bridging the input question and output answer.
@@
  • CoRE is our proposed natural language programming method with LLM as an interpreter.
@@ 4.4 Implementation Details
Our framework and all baselines are implemented by PyTorch, an open-source library. We follow the implementation setting of the OpenAGI platform [14] for Zero-shot and few-shot learnings. We leverage the DSPy framework [23, 24] to apply the CoT strategy to the OpenAGI platform. We also try Program-of-Thought [6] and ReAct [53] strategies on the OpenAGI platform. However, the ReAct strategy requires text observation, which is unsuitable for our OpenAGI task since some observations are in image format, and Program-of-Thought cannot generate executable codes. Thus, we do not include them as the baselines.
@@ 4.5 Experimental Analysis
The experiment results on the OpenAGI benchmark are shown in Table 1. Each row stands for a type of task, each column represents the planning schema of an LLM interpreter, and every four columns are the results of the same LLM interpreter. From the results, we can conclude that our CoRE planning schema is significantly better on average performance than any baseline under both Mixtral and GPT-4 as the interpreters. When using Mixtral as the interpreter, our CoRE system strategy outperforms Zero-shot and CoT under each type of task, and is significantly better than Few-shot learning on Task 2 and average score, though worse on Task 3 and slightly worse on Task 1. When using GPT-4 as the interpreter, CoT, Few-shot has similar performance on Task 1 and Task 3, while on Task 2 and average score, CoRE is still the best. We should note that it is unfair to compare our CoRE with Few-shot learning since we do not directly provide the output format and output example in the prompt. However, under such a situation, our CoRE planning strategy is still better than the Few-shot on average. We also find that using the same CoRE program, the LLM may perform differently when using different LLM interpreters. Thus, the performance of our natural language programming is also determined by the natural language understanding ability of the LLM interpreter.
@@
5 Conclusion and future work
In this study, we introduce a novel system, CoRE, which is used for Code Representation and Execution. CoRE is designed to bridge natural language programming, pseudo-code, and flow programming through the development of a unified CoRE language. Our system leverages Large Language Models (LLMs) as interpreters to process and execute natural language instructions. Throughout execution, the interpreter dynamically retrieves necessary information, utilizes appropriate external tools, and navigates through instructions based on previous outputs. The experimental outcomes validate the efficacy of the CoRE system in natural language programming.
While our initial model demonstrates promising results, it currently relies on manually crafted programs, which may introduce inefficiencies due to the inherent ambiguities of natural language. To address this, future research could explore the development of automated systems for generating natural language programming instructions. This automation would help standardize instruction clarity and precision, potentially improving system performance. Additionally, a future direction is to expand CoRE’s language support to facilitate international use and implement real-time debugging features to aid in education and assist novice programmers, further broadening the system’s utility and accessibility.

### v2

# AIOS Compiler: LLM as Interpreter for Natural Language Programming and Flow Programming of AI Agents

@@ Abstract
Since their inception, programming languages have trended towards greater readability and lower barriers for programmers. Following this trend, natural language can be a promising type of programming language that provides great flexibility and usability and helps towards the democracy of programming. However, the inherent vagueness, ambiguity, and verbosity of natural language pose significant challenges in developing an interpreter that can accurately understand the programming logic and execute instructions written in natural language. Fortunately, recent advancements in Large Language Models (LLMs) have demonstrated remarkable proficiency in interpreting complex natural language. Inspired by this, we develop a novel system for Code Representation and Execution (CoRE), which employs LLM as interpreter to interpret and execute natural language programs (NLPg). The proposed system unifies natural language programming, pseudo-code programming, and flow programming under the same representation for constructing language agents, while LLM serves as the interpreter to interpret and execute the agent programs. In this paper, we begin with defining the programming syntax that structures natural language instructions logically. During the execution, we incorporate external memory to minimize redundancy. Furthermore, we equip the designed interpreter with the capability to invoke external tools, compensating for the limitations of LLM in specialized domains or when accessing real-time information. This work is open-source at https://github.com/agiresearch/CoRE, https://github.com/agiresearch/OpenAGI, and https://github.com/agiresearch/AIOS.
@@ 2.2 Large Language Models and AI Agents for Problem Solving
Large Language Models (LLMs) have emerged as powerful tools for problem solving, encompassing tasks in reasoning, planning, and code generation. LLM reasoning typically involves decomposing a complex task into a sequence of steps, also known as a reasoning chain [49]. Prominent approaches in LLM reasoning include Chain-of-Thought (CoT) and its derivatives [49, 26]. To further improve the reasoning ability of LLM, several work has been proposed. The Self-consistency method [47] samples multiple reasoning paths and selects the most consistent outcome by voting. Additionally, classical data structures like trees and graphs are utilized to enhance reasoning efficiency and accuracy in fewer steps [52, 3]. Apart from reasoning, planning is also an important task that can be used to solve problems. LLM Planning involves generating a series of actions to achieve the predefined goals [16]. Recent advancements include direct prompting of LLMs for planning tasks, showing promising results [20, 45, 11]. Finite state machines have been integrated into LLM to enhance the planning ability [29, 51]. ReAct [53] proposes to leverage external tools like search engine to enhance the LLM planning. Besides, considering the powerful ability of LLM in programming, recent work propose to generate programming code to solve problems [32, 22, 31, 6, 23, 40, 36]. Furthermore, the “self-reflection” mechanism [33, 39, 44] enables LLMs to critique their own outputs, significantly enhancing performance in tasks such as reasoning [3] and code generation [7]. In contrast to existing methods that directly use LLMs for generating solutions, the proposed CoRE system utilizes LLMs as interpreters, executing solutions designed by humans to address complex questions. This approach leverages human creativity in solution design, coupled with LLM’s ability, to enhance problem-solving capabilities in natural language programming contexts.
Figure 3: An overview of the CoRE LLM interpreter system.
@@ 3.1 CoRE Language Syntax
To organize natural language instructions, we define the basic structural representation for each step, which consists of four components. An example can be found in Figure 1.
  • Step Name: Each step in the program is uniquely identified by a step name. This identifier is analogous to function identifiers in traditional programming languages, which facilitates navigation and reference within the program structure, ensuring that each operation within the program can be distinctly addressed and accessed.
  • Step Type: The step type categorizes the nature of the operation being performed in each step, analogous to control structures in conventional programming. We define three primary step types:
@@
    • Decision: Corresponding to conditional statements (e.g., “if-else”), this step involves branching the program flow based on evaluated conditions, leading to multiple potential paths.
    • Terminal: Similar to the “end” or “return” statement, this step marks the conclusion of the program, indicating that no further steps are to be executed.
  • Step Instruction: The step instruction explicates the task to be conducted at a step. This component is integral as it provides the instruction and content for execution, paralleling the statement block in traditional programming languages. By demonstrating operations in natural language, NLPg lowers the barrier to programming, making it more readable for non-expert programmers.
  • Step Connection: Step connections define the progression from one step to another, establishing the flow of the program execution. In process steps, a single subsequent step is specified. In decision steps, multiple pathways are delineated based on conditions. Terminal steps, by definition, do not lead to any future steps, indicating the end of program execution.
For each step in the program, the above four components are separated by “:::” (as illustrated in the CoRE language in Figure 1). Other special tokens can also be used to separate different components.
In programming languages, there are three basic control constructs in programming [9, 41]: sequence, selection and iteration. These three basic constructs can be easily designed within the CoRE language.
  • Sequence: Sequence in programming is the execution of statements in a linear order, with each statement leading to the next. In the CoRE framework, this construct is designed by setting the “Step Connection” to point to the subsequent step. Each step operates under the Process type until the sequence concludes.
@@
  • Iteration: Iteration involves repeating a set of operations until a certain condition is met, akin to loops in conventional programming. In the CoRE framework, we utilize a step with the Decision type to assess whether the loop condition has been fulfilled. At the end of one loop cycle, the “Step Connection” is configured to point back to the previous Decision step, thereby enabling the continuation of the loop.
@@ 3.2 LLM as Interpreter
In this section, we will discuss how the CoRE system utilizes a Large Language Model (LLM) as an interpreter to execute programs written in the CoRE language. We will demonstrate the execution of a single step within the CoRE system, which is illustrated in Figure 3. More specifically, the system executes a single step in four procedures. First of all, the interpreter determines the useful information to execute the current step. Then the interpreter will integrate all relevant information to construct the prompt. Based on the generated prompt, the interpreter will generate response and may utilize tools to execute the current step. Finally, after executing the current step, the interpreter will determine the next step based on step type and execution results. We will introduce the four parts in details.
@@ 3.2.1 Observation Retrieval from Memory
This initial procedure is critical since it sets the stage for the entire execution process of the current step. Figure 4 shows an example. The system’s memory serves as a repository of all prior observations related to the program, where the observation represents the results of tool execution, such as search results. During this phase, the interpreter scans the memory to identify records that are relevant to the current instruction. This selective retrieval ensures that the interpreter’s decisions are informed by accurate and contextually relevant data, which is crucial for the successful execution of the program.
@@ 3.2.2 Input Prompt Construction
Constructing the prompt is essentially about synthesizing the information into a comprehensive and coherent query that the LLM can understand and respond to effectively. This involves combining multiple information into a single, structured prompt that guides the LLM towards generating the most appropriate and contextually relevant response. In the CoRE system, the interpreter constructs a detailed prompt with four elements:
@@ 3.2.3 Output Analysis
While the LLM can generate direct responses, complex tasks may require capabilities beyond its immediate scope. Incorporating the use of specialized tools when necessary extends the LLM’s capabilities, allowing the system to handle a broader range of tasks effectively. A demonstrative example of the execution process is shown in Figure 5. In the CoRE system, the interpreter will make a decision about if or not to employ specialized tools based on the LLM’s initial response and the demands of the task at hand, which ensures that the system remains highly functional and versatile, actively solving problems rather than merely processing the language prompt for the current step. Specifically, if tool usage is warranted, the system will select the suitable tool, configure it with the necessary parameters, execute it, and integrate the output into the ongoing process.
3.2.4 Branching Analysis
@@
Determining the appropriate next step in the program is critical, especially in multi-branch scenarios where different outcomes can lead to different subsequent actions. Figure 6 shows an example. In the CoRE language interpreter, the Decision steps indicate multiple branches with the corresponding conditions. The interpreter uses LLM to decide if the prompt satisfies the natural language described branching condition or not and which next step to take. This adaptive approach allows the system to navigate through decision points effectively, ensuring logical progression toward the program’s goals.
@@ 4.1 Backbone Large Language Model (LLM)
We conduct experiments on both closed-source and open-source LLMs:
  • GPT-4 [37] (Closed-source) is a generative pre-trained transformer of OpenAI. In this work, we use the GPT-4-1106-preview version.
@@ 4.2 Planning Schema of LLMs
  • Chain-of-Thought (CoT) [49] induces the LLM to generate a coherent language sequence that serves as a meaningful intermediate step bridging the input query and the output answer.
@@
  • CoRE is our natural language programming method with LLM as an interpreter.
@@ 4.4 Implementation Details
Our framework and all baselines are implemented by PyTorch, an open-source library. We follow the implementation setting of the OpenAGI platform [14] for Zero-shot and few-shot learnings. We leverage the DSPy framework [24, 25] to apply the CoT strategy to the OpenAGI platform. We also tried Program-of-Thought [6] and ReAct [54] strategies on the OpenAGI platform. However, the ReAct strategy requires text observation, which is unsuitable for our OpenAGI task since some observations are in image format, and Program-of-Thought cannot generate executable codes. Thus, we did not include them as the baselines.
@@ 4.5 Experimental Analysis
The experiment results on the OpenAGI benchmark are shown in Table 1. Each row stands for a type of task, each column represents the planning schema of an LLM interpreter, and every four columns are the results of the same LLM interpreter. From the results, we can see that our CoRE planning schema is better on average performance than any baseline under both Mixtral and GPT-4 as the interpreters. When using Mixtral as the interpreter, CoRE outperforms Zero-shot and CoT under each type of task, and is better than Few-shot learning on Task 2 and average score, though worse on Task 3 and slightly worse on Task 1. When using GPT-4 as the interpreter, CoT, Few-shot has similar performance on Task 1 and Task 3, while on Task 2 and average score, CoRE is still the best. It may be worth noting that it is unfair to compare CoRE with Few-shot learning since we do not directly provide the output format and output example in the prompt. However, even without using such examples, the CoRE planning strategy is still better than the Few-shot strategy on average. We also find that even for the same CoRE program, the system may perform differently when using different LLM as interpreters, which means that the performance of natural language programming depends on the natural language understanding ability of the LLM interpreter.
@@
5 Conclusions and Future Work
In this study, we introduce a novel system, CoRE, for Code Representation and Execution. CoRE is designed to bridge natural language programming, pseudo-code, and flow programming through the development of a unified CoRE language for the construction of AI Agents. CoRE leverages natural language as the programming interface, which lowers the programming barrier and advocates the democracy of programming, so that even ordinary users can create their AI Agents. Our system leverages Large Language Models (LLMs) as interpreters to process and execute natural language instructions. Throughout execution, the interpreter dynamically retrieves necessary information, utilizes appropriate external tools, and navigates through instructions based on previous outputs. The experimental outcomes validate the efficacy of the CoRE system in natural language programming.
While CoRE demonstrates promising results, it currently relies on manually crafted programs, which may introduce inefficiencies due to the inherent ambiguities of natural language. To address this, future research could explore the development of automated systems for generating natural language programming instructions. This automation would help standardize instruction clarity and precision, potentially improving system performance. Additionally, a future direction is to expand CoRE’s language support to facilitate international use and implement real-time debugging features to aid in education and assist novice programmers, further broadening the system’s utility and accessibility.